OpenStudy (anonymous):

in the square root of y minus the square root of 7 equals the square root of y+7, what is y?

OpenStudy (marihelenh):

Is this what it looks like? \[\sqrt{y}-\sqrt{7}=\sqrt{y+7}\]

OpenStudy (marihelenh):

If it is, you would square both sides to get rid of the square roots. Then you could solve for y.

OpenStudy (mathstudent55):

Did you get a solution?

OpenStudy (anonymous):

I just needed to know whether the solution contained any roots

OpenStudy (mathstudent55):

You need to solve it to find out.